### v1

# DAUNet: A Lightweight UNet Variant with Deformable Convolutions and Parameter-Free Attention for Medical Image Segmentation

Abstract
Medical image segmentation plays a pivotal role in automated diagnostic and treatment planning systems. In this work, we present DAUNet, a novel lightweight UNet variant that integrates Deformable V2 Convolutions and Parameter-Free Attention (SimAM) to improve spatial adaptability and context-aware feature fusion without increasing model complexity. DAUNet’s bottleneck employs dynamic deformable kernels to handle geometric variations, while the decoder and skip pathways are enhanced using SimAM attention modules for saliency-aware refinement. Extensive evaluations on two challenging datasets—FH-PS-AoP (fetal head and pubic symphysis ultrasound) and FUMPE (CT-based pulmonary embolism detection)—demonstrate that DAUNet outperforms state-of-the-art models in Dice score, HD95, and ASD, while maintaining superior parameter efficiency. Ablation studies highlight the individual contributions of deformable convolutions and SimAM attention. DAUNet’s robustness to missing context and low-contrast regions establishes its suitability for deployment in real-time and resource-constrained clinical environments.
{IEEEkeywords}
Deformable convolutions, Medical image segmentation, Parameter-free attention, Pulmonary embolism detection, Ultrasound imaging, UNet variant.
1 Introduction
Medical image segmentation is a foundational task in computer-assisted diagnosis, enabling the precise localization and delineation of anatomical structures that are critical for clinical interpretation, surgical planning, and disease monitoring. Accurate and automated segmentation reduces manual effort and inter-observer variability, particularly in high-throughput clinical settings. Despite significant advances achieved through convolutional neural networks (CNNs), especially the widely adopted UNet architecture [1], key challenges persist, most notably in achieving robustness, handling anatomical variability, and maintaining computational efficiency.
Although effective in many scenarios, the classical UNet architecture presents several limitations. Its use of fixed-grid convolutions restricts adaptability to variable-sized features and irregular organ boundaries. Additionally, UNet often struggles in low-contrast or noisy environments, common in modalities such as ultrasound [2, 3] and CT angiography, where anatomical boundaries are not clearly visible [4, 5]. Moreover, UNet lacks mechanisms to capture long-range dependencies, which are crucial for modeling global context in complex medical images.
To overcome these shortcomings, recent works have explored enhancements to UNet via transformer-based modules and attention mechanisms [6, 7]. For instance, Masoudi et al. [6] proposed FAT-Net, which augments a UNet-style backbone with transformer branches to capture long-range interactions and feature adaptation modules to suppress background noise. Similarly, Zhang et al. [8] introduced TransAttUNet, incorporating a Self-Aware Attention (SAA) module that integrates Transformer Self-Attention (TSA) and Global Spatial Attention (GSA) to improve multi-scale feature fusion. Other methods such as DSEUNet [9] and MISSFormer [10] attempt to bridge CNN and transformer paradigms. DSEUNet deepens the UNet backbone while introducing Squeeze-and-Excitation (SE) blocks [11] and hierarchical supervision. MISSFormer, on the other hand, employs enhanced transformer blocks and multi-scale fusion to balance local and global feature representation.
General-purpose models such as MedSAM [12], adapted from the Segment Anything Model (SAM), offer prompt-based segmentation across various modalities. Trained on over 1.5 million image–mask pairs, MedSAM shows strong performance on CT, MRI, and endoscopy images. However, limitations remain due to the underrepresentation of certain modalities (e.g., mammography) and imprecise vessel boundary segmentation when using bounding-box prompts.
While the aforementioned models demonstrate commendable segmentation performance, they often suffer from high computational complexity and slower inference, limiting their suitability for real-time or resource-constrained environments.
Hybrid models like H2Former [13] and SCUNet++ [14] further aim to unify the strengths of CNNs and transformers. H2Former leverages hierarchical token-wise and channel-wise attention to model both local and global dependencies. SCUNet++ integrates CNN bottlenecks and dense skip connections to improve pulmonary embolism (PE) segmentation. Although SCUNet++ achieves high Dice scores on PE datasets, it tends to produce blocky segmentation outputs on large lesions and has a substantial parameter burden. Other methods, such as CE-Net [15], augment UNet with Dense Atrous Convolution (DAC) and Residual Multi-kernel Pooling (RMP) blocks to improve feature representation, but their multi-branch architectures increase memory requirements and limit scalability.
Motivated by the need for efficient, adaptable, and robust segmentation models suitable for real-world clinical deployment, we propose DAUNet, a lightweight and effective UNet-based architecture featuring two key innovations:
  • Improved Bottleneck: A lightweight deformable convolution-based bottleneck module [16, 17] that introduces dynamic, spatially adaptive receptive fields. This design enables the model to better capture geometric deformations and irregular anatomical boundaries.
  • Improved Decoder: A parameter-free attention mechanism (SimAM) [18] is integrated into the decoder and skips connections to enhance spatial feature representation and facilitate efficient feature fusion, without increasing model complexity.
To demonstrate its effectiveness, we evaluate DAUNet on two challenging medical image segmentation tasks: (1) fetal head and pubic symphysis segmentation from transperineal ultrasound using the FH-PS-AoP dataset [19], and (2) pulmonary embolism detection in CT angiography using the FUMPE dataset [20]. Both tasks are characterized by substantial anatomical variability, low-contrast regions, and limited contextual information, factors that commonly impair the performance of conventional models. Through comprehensive experiments, DAUNet achieves superior segmentation accuracy, robustness to missing context, and significantly improved parameter efficiency compared to state-of-the-art methods. These results, supported by ablation and robustness analyses, highlight DAUNet’s practical potential for deployment in real-time and resource-constrained clinical environments.
The remainder of this paper is organized as follows: Section 2 describes the proposed DAUNet architecture in detail. Section 3 outlines the experimental setup, datasets, and evaluation metrics. In Section 4, we present and discuss quantitative and qualitative results, including ablation studies and robustness analysis. In Section 5, we provides a discussion on clinical impact. Finally, Section 6 concludes the paper and outlines directions for future work.
2 Methodology
The proposed framework is built upon the UNet architecture with two key modifications: Deformable V2 Convolutions [16, 17] in the bottleneck block and Parameter-Free Attention [18] in the decoder block.
2.1 Deformable Convolution
Deformable convolution [16, 17] extends the conventional convolution operation by incorporating spatial offsets that adapt dynamically based on input features. This mechanism enables the network to focus on pertinent regions, effectively capturing spatial deformations, such as scaling, rotation, and complex anatomical structures.
Figure 1 illustrates the operational differences between standard and deformable convolutions. The left panel depicts a standard $3\times 3$ convolution, where sampling locations are fixed. In contrast, the right panel shows a deformable convolution, where each sampling location is adjusted by a learnable offset, depicted by the black arrows.
Figure 1: Comparison between standard and deformable convolutions. (a) Standard convolution with fixed sampling grid. (b) Deformable convolution with learnable offsets. The offsets are dynamically adjusted based on input features, allowing the model to better capture spatial variations.
2.1.1 Deformable Convolution Operation
For a given input feature map $\mathbf{F}$ of size $C\times H\times W$ and a convolution kernel $\mathbf{K}$ of size $C\times w\times w$, the output at location $\mathbf{p}$ is computed as:
$Y(\mathbf{p})=\sum_{\mathbf{k}\in\mathcal{R}}\mathbf{K}(\mathbf{k})\cdot\mathbf{F}(\mathbf{p}+\mathbf{k}+\Delta\mathbf{p_{k}}),$ (1)
where $\mathcal{R}$ is the receptive field, and $\Delta\mathbf{p_{k}}$ is the learnable offset. Since $\mathbf{p}+\mathbf{k}+\Delta\mathbf{p_{k}}$ may be a non-integer location, bilinear interpolation is applied to approximate the feature values at fractional coordinates.
2.1.2 Modulated Deformable Convolution
To further enhance flexibility, a modulation scalar $\alpha_{k}\in[0,1]$ is introduced [17], allowing the network to selectively emphasize or suppress specific regions:
$Y(\mathbf{p})=\sum_{\mathbf{k}\in\mathcal{R}}\alpha_{k}\cdot\mathbf{K}(\mathbf{k})\cdot\mathbf{F}(\mathbf{p}+\mathbf{k}+\Delta\mathbf{p_{k}}).$ (2)
The modulation scalars are generated through an auxiliary convolutional layer followed by a sigmoid activation. This mechanism enables the network to focus more effectively on salient features, further enhancing the segmentation accuracy.
2.2 Parameter-Free Attention: SimAM
Figure 2: Schematic illustration of the SimAM attention mechanism. For each neuron, an energy-based evaluation is performed using surrounding spatial context, followed by an attention weighting operation. The resulting map highlights informative regions without introducing learnable parameters.
To enhance feature representation without increasing model complexity, we incorporate the Simple Attention Module (SimAM) [18] into the decoder blocks of our architecture. SimAM is a parameter-free attention mechanism inspired by neuroscience theories, particularly the spatial suppression theory, which suggests that neurons with higher activation energies are less informative.
As illustrated in Figure 2, SimAM performs attention modulation by evaluating each neuron’s importance through an energy-based function, which considers both the neuron’s deviation from the mean and the dispersion of surrounding neurons. This formulation allows SimAM to assign lower weights to high-energy (i.e., less informative) activations while highlighting those contributing meaningful context.
Given a feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, SimAM computes an attention weight for each neuron based on the following energy function:
$E_{t}=(x_{t}-\mu_{t})^{2}+\lambda\sum_{i\neq t}(x_{i}-\mu_{t})^{2},$ (3)
where $\mu_{t}$ is the mean of all neurons in the same channel excluding $x_{t}$, and $\lambda$ is a hyperparameter controlling the importance of surrounding neurons and it is set to be $\lambda=1\times 10^{-4}$.
The attention weight $a_{t}$ for each neuron is computed as:
$a_{t}=\sigma\left(\frac{1}{E_{t}+\epsilon}\right),$ (4)
where $\sigma(\cdot)$ denotes the sigmoid activation, and $\epsilon$ is a small constant to avoid division by zero.
The refined output $\mathbf{X}^{\prime}$ is obtained by element-wise multiplication of the original feature map and the attention weights:
$\mathbf{X}^{\prime}=\mathbf{X}\odot\mathbf{A},$ (5)
where $\odot$ denotes element-wise multiplication and $\mathbf{A}$ is the attention map composed of $a_{t}$ values across all spatial positions.
By applying SimAM, our model selectively amplifies informative features while suppressing less relevant ones—all without adding any additional parameters—thus achieving both enhanced accuracy and computational efficiency.
2.3 Proposed Architecture
Figure 3: Overview of the proposed DAUNet architecture. The network is a lightweight variant of the UNet, incorporating key modifications in the bottleneck and skip connection paths. The bottleneck block is redesigned with a sequence of three operations: a $1\times 1$ convolution for channel compression, a $3\times 3$ Deformable Convolution V2 for adaptive spatial modeling, followed by another $1\times 1$ convolution and a SimAM attention block. Additionally, the skip connections are augmented with SimAM modules to enhance feature fusion and emphasize informative activations without increasing parameter count.
The architecture of the proposed DAUNet is shown in Figure 3. It retains the encoder–decoder structure of the classical UNet but introduces key improvements to enhance representational power and efficiency. In both the encoder and the decoder blocks, all convolution operations uses a $3\times 3$ size filters. The bottleneck block is significantly redesigned by replacing the conventional convolution with a compound structure composed of four consecutive operations. First, a $1\times 1$ convolution is used to compress the input channels to a quarter of the target output channels, and control computational cost. This is followed by a $3\times 3$ Deformable Convolution V2 layer [16, 17], which introduces spatial adaptability by learning dynamic offsets to the sampling grid, effectively capturing geometric variations in the input features. A second $1\times 1$ convolution is then applied to project the features back to the original channel dimension. Finally, a SimAM module is appended to refine the output by emphasizing spatially informative activations based on an energy-based criterion, without introducing any learnable parameters.
In addition to the modified bottleneck, we also enhance all skip connection pathways by integrating SimAM modules before merging the encoder and decoder features. This modification helps suppress irrelevant or redundant activations and strengthens the transfer of semantically rich features across the network. These lightweight yet effective modifications improve segmentation accuracy and generalization, especially in challenging medical imaging scenarios, without significantly increasing the model’s parameter count or computational burden.
3 Experimental Setup
The experimental setup of the study is defined as follows:
3.1 Training Setup
The CNN models were trained using the Adam optimizer [21], with a fixed learning rate of $0.0001$. Training was performed on an NVIDIA RTX A4500 GPU, running each model for 150 epochs with a batch size of $12$. To maintain consistency during evaluation, all input images were resized to a fixed resolution of $256\times 256$ pixels.
The models were developed using the PyTorch framework [22], leveraging GPU acceleration through CUDA version 12.2. For the proposed approach, a hybrid loss function was employed, combining Dice loss with Weighted Binary Cross-Entropy (BCE) [23] to effectively guide the learning process.
We also used geometric transformation-based data augmentation methods such as adaptive zooming, rotation, and random flipping. By boosting the variety of rare class appearances, these augmentations not only add diversity to the training data to avoid overfitting but also aid in addressing class imbalance.
3.2 Dataset
The proposed model is evaluated on two datasets:
  • Pubic Symphysis and Fetal Head Detection (FH-PS-AoP): The FH-PS-AoP dataset [19] comes from the Pubic Symphysis–Fetal Head Segmentation Challenge, which deals with transperineal ultrasound imaging. It includes 2D B-mode ultrasound images from different regions of China, 85 gathered with standard ultrasound machines. The dataset contains information about pregnant women between 18 and 46 years old. Annotation was conducted using Pair software and was performed by a group of seven annotators composed of two senior specialists and five undergraduates. The entire dataset contains 4,000 images for training and 700 for testing. For model training, the 4,000 training samples were divided into an $80:20$ ratio for training and validation to monitor potential overfitting. Lastly, the 700 test samples were used for performance evaluation.
  • Pulmonary Embolism Detection (FUMPE Dataset): The FUMPE dataset (Ferdowsi University of Mashhad’s Pulmonary Embolism dataset)  [20] contains the three-dimensional computed tomography angiography (CTA) scans of 35 patients, amounting to 8792 image slices. Two radiologists manually reviewed all the CTA scans using a semi-automated segmentation method to derive reference labels. In total, the dataset consists of 3,438 PE (pulmonary embolism) annotated regions. Notably, approximately 67% of these regions occur in the peripheral pulmonary arteries, which positions this dataset to be highly advantageous for prototyping and benchmarking sophisticated CAD systems. After a comprehensive review, we completed the collection of CTPA images for 35 patients, totaling 8792 CTPA images, along with updated annotations by the doctors. For performance evaluation, the dataset was divided into an $80:20$ ratio for training and validation. And performance metric were reported on unseen validation samples as suggested in [14].
3.3 Evaluation Metrics
To quantitatively assess the segmentation performance of DAUNet and compare it against other state-of-the-art models, we employ three widely adopted metrics in medical image segmentation: Dice Similarity Coefficient (DSC) [24], 95th percentile Hausdorff Distance (HD95) [25], and Average Symmetric Surface Distance (ASD) [26]. These metrics provide a comprehensive evaluation across both region-based and boundary-based performance criteria. Lastly, we use parameter count as a metric for model complexity.
3.3.1 Dice Similarity Coefficient (DSC)
The Dice Similarity Coefficient measures the overlap between the predicted segmentation mask $P$ and the ground truth mask $G$. It is defined as:
$\text{DSC}=\frac{2|P\cap G|}{|P|+|G|},$ (6)
where $|\cdot|$ denotes the cardinality of a set. DSC ranges from 0 to 1, where a higher value indicates better agreement.
3.3.2 Hausdorff Distance (HD95)
Hausdorff Distance quantifies the maximum boundary deviation between the predicted and ground truth masks. We use the 95th percentile version for robustness to outliers:
$\begin{split}\text{HD}_{95}(P,G)=\max\Big\{\sup_{p\in\partial P}\inf_{g\in\partial G}\|p-g\|,\\ \sup_{g\in\partial G}\inf_{p\in\partial P}\|g-p\|\Big\}_{95\%},\end{split}$ (7)
where $\partial P$ and $\partial G$ are the boundaries of the predicted and ground truth masks. Lower HD95 values indicate better boundary alignment.
3.3.3 Average Symmetric Surface Distance (ASD)
ASD measures the average bidirectional distance between the boundaries of prediction and ground truth:
$\begin{split}\text{ASD}(P,G)=\frac{1}{|\partial P|+|\partial G|}\Big(\sum_{p\in\partial P}\min_{g\in\partial G}\|p-g\|\\ +\sum_{g\in\partial G}\min_{p\in\partial P}\|g-p\|\Big).\end{split}$ (8)
ASD provides an interpretable measure of the mean boundary error, with lower values indicating better segmentation precision.
3.3.4 Model Complexity (Parameter Count)
In addition to accuracy-based metrics, we also report the number of trainable parameters as an indicator of model complexity. Lower parameter counts are desirable in clinical scenarios involving real-time processing or deployment on edge devices.
4 Experimental Results
We conduct extensive experiments on two challenging medical image segmentation tasks, including pubic symphysis and fetal head segmentation from ultrasound (FH-PS-AoP dataset), and pulmonary embolism detection from CT angiography scans (FUMPE dataset), to evaluate the performance of the proposed DAUNet model.
4.1 Comparison with State-of-the-Art Methods
Table 1: Quantitative comparison of DAUNet and state-of-the-art models on the FH-PS-AoP dataset. Performance is reported using Dice Similarity Coefficient (DSC $\uparrow$), 95th percentile Hausdorff Distance (HD95 $\downarrow$), and Average Symmetric Surface Distance (ASD $\downarrow$). Bold values indicate the best results.
Methods DSC $\uparrow$ HD95 $\downarrow$ ASD $\downarrow$ Param (M) $\downarrow$
PSFH FH PS PSFH FH PS PSFH FH PS
UNet [1] 80.22 81.42 79.02 15.87 17.42 14.33 4.88 5.09 4.67 31.03
Attention-UNet [27] 85.55 85.86 85.24 15.66 17.48 13.85 4.83 4.92 4.75 34.88
TransUNet [28] 87.34 93.02 81.66 12.30 13.25 11.36 3.67 3.58 3.76 105.28
SwinUNet [29] 87.45 92.06 82.84 12.50 13.96 11.05 3.74 3.44 4.04 27.17
SCUNet++ [14] 82.90 88.40 77.39 13.96 18.76 9.15 5.60 5.30 5.90 60.11
DSEUNet [9] 88.65 92.44 84.86 14.57 16.47 12.67 3.9 4.46 3.34 62.9
UPerNet [30] 86.45 92.24 80.67 13.47 14.36 12.59 4.11 4.43 3.80 31.3
MissFormer [10] 86.09 91.58 80.61 14.71 15.78 13.65 4.38 4.81 3.96 42.3
H2Former [13] 88.98 92.64 85.32 12.57 14.06 11.09 3.57 4.19 2.96 33.9
CE-Net [15] 88.79 92.68 84.9 14.03 14.57 13.50 3.79 3.82 3.76 29.0
MedSAM [12] 87.22 90.59 83.85 12.93 16.36 9.51 4.02 5.26 2.78 93.7
TransAttUNet [8] 85.79 91.11 80.47 12.46 15.66 9.27 4.36 5.53 3.19 25.96
FAT-Net [6] 88.35 92.65 84.05 13.28 14.12 12.45 4.28 3.99 4.57 30.00
DAUNet (Proposed) 89.09 92.87 85.31 10.37 12.76 7.98 3.70 4.44 2.95 20.47
Table 2: Performance comparison of DAUNet and competing models on the FUMPE dataset. DAUNet achieves the best segmentation accuracy (DSC $\uparrow$) and lowest boundary error (HD95 $\downarrow$) with fewer parameters, demonstrating superior efficiency and robustness.
Methods DSC $\uparrow$ HD95 $\downarrow$ Param (M) $\downarrow$
UNet [1] 77.91 6.86 31.03
Unet ++ [31] 77.16 5.80 34.96
SwinUNet [29] 60.80 20.20 25.91
ResD-UNet [32] 76.48 22.25 50.83
TransAttUNet [8] 82.82 3.2 25.96
SCUNet++ [14] 83.47 3.83 60.11
FAT-Net [6] 84.44 3.67 30.00
DAUNet (Proposed) 88.80 2.57 20.47
Tables 1 and 2 present a quantitative comparison between DAUNet and several state-of-the-art segmentation models, including UNet, Attention-UNet, TransUNet, SwinUNet, SCUNet++, TransAttUNet, and FAT-Net. Performance is evaluated using three key metrics: Dice Similarity Coefficient (DSC), 95th percentile Hausdorff Distance (HD95), and Average Symmetric Surface Distance (ASD).
For the FH-PS-AoP dataset, Table 1 presents a detailed comparison of DAUNet with recent state-of-the-art models. DAUNet achieves the highest average Dice Similarity Coefficient (DSC) on the PSFH class (89.09%) and ranks second-best for the FH (92.87%) and PS (85.31%) classes—trailing TransUNet [28] and H2Former [13], respectively, by small margins. It also delivers the lowest Hausdorff Distance (HD95) across all classes and achieves top boundary accuracy (ASD) for PS, while ranking third-best for PSFH (3.70 vs. 3.67 for TransUNet [28] and 3.57 for H2Former [13]) and FH (4.44 which is comparable).
Despite these slight trade-offs, DAUNet excels in efficiency—requiring only 20.47M parameters—substantially fewer than heavier models like TransUNet [28] (105.28M) and MedSAM [12] (93.7M), while matching or outperforming them in all major metrics. Compared to the baseline UNet [1], it improves the mean DSC by nearly 9%, and significantly reduces boundary errors (HD95 and ASD) by over 5 units.
Even against compact models like TransAttUNet [8] (25.96M) and SwinUNet [29] (27.17M), DAUNet achieves superior accuracy with fewer parameters. This balance between accuracy and model compactness makes it especially well-suited for real-time and edge-based ultrasound deployments [33].
Similarly, on the FUMPE dataset, proposed DAUNet achieves the best performance with a Dice score of 88.80% and HD95 of 2.57, while maintaining a significantly lower parameter count (20.47M) compared to most competing methods. Notably, even models with over 60M parameters (e.g., SCUNet++) fall short of matching the segmentation accuracy and robustness achieved by DAUNet. It is worth noting that our model was evaluated on datasets with two different segmentation input sizes, consistently demonstrating strong performance across both smaller and larger lesion areas.
4.2 Qualitative Analysis
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Ground-Truth UNet SCUNet++ TransAttUNet FAT-Net DAUNet (Proposed)
Figure 4: Segmentation results of different models on FH-PS-AoP dataset: (a) Input image, (b), The ground truth mask, (c) UNet, (d) SCUNet++, (e) TransAttUNet (f) FAT-Net , (g) DAUNet (Proposed). The contours around the prediction masks are the ground truth mask contours.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Ground-Truth UNet SCUNet++ TransAttUNet FAT-Net DAUNet (Proposed)
Figure 5: Segmentation results of different models on FUMPE dataset: (a) Input image, (b), The ground truth mask, (c) UNet, (d) SCUNet++, (e) TransAttUNet (f) FAT-Net , (g) DAUNet (Proposed).
Figure 4 and Figure 7 provide visual comparisons of segmentation masks generated by DAUNet and other competing methods on representative samples from both datasets. As illustrated in Figure 4, DAUNet produces more accurate and smoother boundary delineations, closely aligning with the ground truth annotations, even in challenging low-contrast regions or under partial occlusion. A similar performance can be observed in Figure 7, where proposed model accurately identify the pulmonary embolism regions. It is worth nothing that the proposed model achieve best performance in two different imaging modalities, consisting of different sized regions of interest and number of classes demonstrating a good generalization.
4.3 Parameter Efficiency vs. Performance
Figure 6: Trade-off between model complexity and segmentation accuracy on the PSFH subset of the FH-PS-AoP dataset. The x-axis represents the number of trainable parameters (log scale), while the y-axis shows the Dice Similarity Coefficient (DSC). DAUNet (Proposed) achieves the highest DSC with the fewest parameters among all methods, demonstrating its superior efficiency and accuracy.
To highlight the efficiency of DAUNet, Figure 6 illustrates the trade-off between model complexity (in terms of parameter count) and segmentation performance (DSC) on FH-PS-AoP dataset. While most transformer-based and heavily parameterized CNN models achieve high accuracy at the cost of computational load, DAUNet attains state-of-the-art performance with significantly fewer parameters. This favorable balance makes DAUNet highly suitable for real-time deployment in clinical workflows and on resource-constrained platforms such as edge devices and mobile ultrasound systems [33].
4.4 Robustness to Missing Context
Input
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
UNet (Offsets)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
UNet (Masks)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DAUNet (Offsets)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
DAUNet (Masks)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Robustness to missing context: Comparison of offset maps and segmentation masks generated by UNet and DAUNet for inputs with cropped regions. The first column shows the original image and output; subsequent columns show results for inputs with missing top-left, top-right, bottom-left, and bottom-right patches. DAUNet demonstrates stronger contextual inference and better preservation of structural boundaries despite incomplete input.
In practical clinical settings, medical images are often incomplete or degraded due to occlusions, artifacts, or limited fields-of-view. Therefore, a reliable segmentation model must be robust to missing spatial context and capable of inferring anatomical structures from partially observed inputs. To evaluate this critical property, we conducted a masking experiment using the FH-PS-AoP dataset. In this experiment, we systematically removed one quadrant from the input image, top-left, top-right, bottom-left, or bottom-right, and evaluated the performance of DAUNet compared to the baseline UNet.
Figure 7 presents a comprehensive visual comparison of the predicted offset maps and segmentation masks under each masking condition. The first column shows the original (unmasked) input along with its corresponding predictions, while subsequent columns depict the results for each of the masked quadrants.
DAUNet demonstrates markedly higher resilience to missing context compared to UNet. Its predicted offset maps remain dense and structured, with vectors that preserve anatomical directionality even when significant regions of the input are absent. This behavior reflects DAUNet’s ability to infer context from the remaining visual cues. In contrast, UNet exhibits sparse or disoriented offsets in the masked scenarios, indicating reduced spatial awareness and compromised localization.
To better understand this phenomenon, we analyzed the receptive fields of both models. For a representative pixel, we visualized its corresponding receptive area contributing to the output. In UNet, the receptive field is fixed and grid-constrained (refer to second row of Figure 7), making it sensitive to occlusions. On the other hand, DAUNet leverages deformable convolutions to dynamically adjust its receptive field based on the visible content. This adaptability is evident in the red and white arrows in the fourth column of Figure 7, where DAUNet modifies its offset patterns to account for the masked input.
The segmentation masks in third and fourth rows of Figure 7, further substantiate these findings. DAUNet consistently produces anatomically plausible segmentations of both the fetal head (green) and pubic symphysis (red), with minimal degradation even under 25% missing input. The output contours remain smooth, accurate, and well-aligned with ground truth boundaries. In contrast, UNet’s performance deteriorates noticeably, with fragmented or distorted segmentations, especially around the fetal head, as indicated by white arrows in the figure.
These results highlight the efficacy of DAUNet’s architectural innovations. The combination of deformable convolution and SimAM attention allows the model to effectively reason over the visible context and compensate for spatial omissions. This robustness to incomplete inputs makes DAUNet well-suited for deployment in real-world medical environments, where noise, occlusions, and partial data are common challenges.
4.5 Ablation Study
To assess the individual contributions of the core components in the DAUNet architecture, we conduct a comprehensive ablation study on the FH-PS-AoP dataset. The goal is to quantify the performance impact of (i) the Deformable V2 Convolution-based bottleneck and (ii) the SimAM parameter-free attention mechanism integrated into the decoder and skip connections.
4.5.1 Impact of Deformable V2 Convolution-Based Bottleneck
Replacing the conventional convolutional layers in the bottleneck with Deformable V2 Convolutions substantially enhances the model’s capacity to adapt to complex anatomical variations. By enabling spatially dynamic receptive fields, this modification allows DAUNet to better localize boundaries in challenging regions. As shown in Table 3, this change alone results in a 7.45% absolute improvement in Dice score (from 80.22% to 87.67%), while simultaneously reducing the parameter count from 31.03M to 20.47M, highlighting its dual benefit of accuracy and efficiency.
4.5.2 Impact of Parameter-Free Attention (SimAM)
Incorporating SimAM modules into the decoder blocks and skip connections improves the model’s spatial feature representation by enhancing the saliency of relevant regions without increasing the parameter count. Although SimAM does not introduce additional weights, its removal results in a noticeable performance drop (Dice: 89.09% to 87.67%, HD95: 10.37 to 11.87), confirming its critical role in refining spatial attention and contextual understanding.
4.5.3 Combined Contribution Analysis
Table 3 summarizes the ablation configurations. Each setup builds progressively upon the baseline UNet model. When both the Deformable V2 bottleneck and SimAM are included, the resulting configuration yields the best performance across all metrics: Dice score improves by +8.87%, HD95 is reduced by 5.50 units, and ASD decreases by 1.18. Furthermore, this performance gain is achieved with a significantly leaner architecture, reducing the total number of parameters by 10.56 million. These findings validate that each component plays a complementary and synergistic role in boosting segmentation accuracy, robustness, and efficiency.
Table 3: Ablation study on FH-PS-AoP dataset. Performance is reported using Dice Similarity Coefficient (DSC), 95th percentile Hausdorff Distance (HD95), and Average Symmetric Surface Distance (ASD).
Proposed Bottleneck SimAM Dice Score (DSC) $\uparrow$ HD95 $\downarrow$ ADS $\downarrow$ Param (M) $\downarrow$
80.22 15.87 4.88 31.03
82.16 14.73 4.92 31.03
87.67 11.87 4.09 20.47
89.09 (+8.87) 10.37 (-5.50) 3.70 (-1.18) 20.47 (-10.56)
5 Discussion
The experimental results confirm that DAUNet achieves state-of-the-art performance on both ultrasound (FH-PS-AoP) and CT (FUMPE) datasets, surpassing classical and transformer-based architectures in both segmentation accuracy and parameter efficiency. The superior Dice scores and reduced HD95 and ASD values demonstrate DAUNet’s ability to precisely capture fine anatomical boundaries, even in low-contrast or noisy conditions typical of real-world clinical data.
Compared to heavily parameterized models such as TransUNet and SCUNet++, DAUNet achieves competitive or superior performance with fewer parameters, making it ideal for deployment in real-time clinical environments and edge devices. This is especially relevant for resource-constrained settings where diagnostic imaging tools are limited, and computational infrastructure is minimal.
5.1 Clinical Impact and Relevance
In clinical scenarios, accurate segmentation of fetal head and pubic symphysis in transperineal ultrasound is essential for assessing fetal positioning and guiding delivery decisions [34]. Errors in segmentation can lead to misjudgment in head station evaluation, potentially increasing the risk of obstructed labor [35]. DAUNet’s robust performance on the FH-PS-AoP dataset highlights its suitability for integration into real-time obstetric ultrasound systems, potentially assisting clinicians in low-resource or emergency settings.
Similarly, the FUMPE dataset targets the detection of pulmonary embolism (PE), a life-threatening condition that demands prompt diagnosis [36]. DAUNet demonstrated excellent boundary precision and robustness in segmenting emboli within peripheral pulmonary arteries, which are often missed by conventional methods. By providing accurate and fast segmentation, DAUNet could be integrated into clinical decision support systems (CDSS) or computer-aided diagnosis (CAD) tools to assist radiologists in detecting PE more efficiently.
Moreover, the architectural design of DAUNet—featuring deformable convolutions and parameter-free attention—ensures that the model generalizes well across different imaging modalities and anatomical structures. This adaptability is clinically valuable, as it reduces the need for extensive retraining and fine-tuning for other challenging applications, such as tumor boundary delineation in oncology or vessel segmentation in cardiology.
From a translational perspective, DAUNet’s low parameter count and fast inference speed make it an excellent candidate for deployment on portable ultrasound machines and embedded systems, thereby enabling point-of-care diagnostics and remote healthcare delivery. This aligns with current trends in mobile health (mHealth), where AI models are increasingly used to bring expert-level diagnostics to underserved populations [37].
6 Conclusion
In this paper, we introduced DAUNet, a lightweight and efficient UNet-based segmentation architecture that integrates Deformable V2 Convolutions and SimAM, a parameter-free attention mechanism, to enhance spatial adaptability and feature selectivity without increasing model complexity. Through extensive evaluation on two challenging datasets, including FH-PS-AoP for ultrasound-based fetal head and pubic symphysis segmentation, and FUMPE for pulmonary embolism detection from CT, DAUNet demonstrated superior performance in terms of segmentation accuracy, boundary preservation, and parameter efficiency compared to existing state-of-the-art models. Ablation studies confirmed the individual and combined contributions of the proposed components, while robustness analysis under missing context scenarios highlighted the model’s ability to infer anatomical structure even with incomplete input. These results underscore DAUNet’s potential for deployment in real-world clinical environments, particularly where computational resources are limited or image quality is compromised. Future work will explore extending the framework to multi-modal and 3D imaging, improving cross-domain generalization through domain adaptation, and further optimizing the model for real-time inference on edge devices.
Acknowledgment
Shujaat Khan acknowledges the support from the King Fahd University of Petroleum & Minerals (KFUPM) under Early Career Research Grant no. EC241027. Additionally, the support from KFUPM Ibn Battuta Global Scholarship Program Grant No: ISP241-COE-872 is gratefully acknowledged.